PATGEN(1) PATGEN(1) NAME patgen-generatepatternsforTeXhyphenation SYNOPSIS patgendictionaryfilepatternfileoutputfiletranslatefile DESCRIPTION Thepatgenprogramreadsthedictionaryfileandproducesthepatternfile,writingatranscriptof itsactionsontheoutputfile.Thedictionaryfileshouldcontainalistofhyphenatedwordsand previously-generatedpatterns. Thepatternsthusgeneratedcanbereadbyinitexforuseinhyphenatingwords.Fora(very)long exampleofpatgen’soutput,see /usr/TeX/lib/tex/macros/hyphen.tex,whichcontainsthepatternsTeXusesforEnglish. Allfilenamesmustbecomplete.Noaddingofdefaultextensionsorpathsearchingisdone. FILES /usr/TeX/lib/tex/macros/hyphen.texPatternsforEnglish. SEEALSO FrankLiang,Wordhy-phen-a-tionbycom-puter,StanfordPh.D.thesis AUTHORS FrankLiangwrotethefirstversionofthisprogram.PeterBreitenlohnermadeasubstantialrevisionin1991forTeX3.ThefirstversionwaspublishedastheappendixtotheTeXwaretechnical report,availablefromtheTeXUsersGroup.HowardTrickeyoriginallyportedittoUnix. 21Mar92 1